OpenStudy (anonymous):

On the graph of f(x) = 8*sin(3pi*x), points P and Q are at consecutive lowest and highest points with P occurring before Q. Find the slope of the line which passes through P and Q. Slope = ? I have tried 48/pi and it didn't work

OpenStudy (anonymous):

ymax-ymin=8-(-8)=16 xmax-xmin=1/6-(-1/6)=1/3 (ymax-ymin)/(xmax-xmin)=48